OpenStudy (anonymous):

intergal of 3x^5 sin(x^3)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}3x^5 \sin(x^3) dx\]

OpenStudy (anonymous):

i chose u= sin(x^3) dv= 3x^5

OpenStudy (anonymous):

you should : let t = x^3 --> dt = 3x^2 dx and \(x^5 = x^2x^3\) \[\rightarrow \int t~sin~t~ dt\]

OpenStudy (anonymous):

now, take it by parts by letting u =t and dv = sint dt

OpenStudy (anonymous):

i got du= 3x^2 cos(x^3) dx

OpenStudy (anonymous):

v= 1/2 x^6

OpenStudy (anonymous):

you got stuck when the exponent of x is bigger and bigger if you go on that way

OpenStudy (anonymous):

oh no wonder it keep going forever

OpenStudy (anonymous):

Try my way!! :)

OpenStudy (anonymous):

happend to 3x^5?

OpenStudy (anonymous):

\(x^5 =x^{3+2}=x^3*x^2\)

OpenStudy (anonymous):

oooooooooooooooo

OpenStudy (anonymous):

what does oooooooooooooooo mean?

OpenStudy (anonymous):

'oh' lol

OpenStudy (anonymous):

got it now, right? :)

OpenStudy (anonymous):

let me try :D

OpenStudy (anonymous):

so for 3 i bring it out side?

OpenStudy (anonymous):

nope, let t = x^3 --> dt = 3x^2 dx "3" is used here--> no need to move it out

OpenStudy (anonymous):